If a business man knows that he will need $150000 in 5 years time, how much does he need to invest now id he has the opportunity to invest at 6.75% compounded quarterly?
Answer by stanbon(75887)

If a business man knows that he will need $150000 in 5 years time, how much does he need to invest now id he has the opportunity to invest at 6.75% compounded quarterly?
-----------
A(t) = P(1+(r/n))^(nt)
-----
150,000 = P(1+(0.0675/4))^(4*5)
-------
150,000 = P(1+0.016875)^20
----
150,000 = P(1.3975)
P = $107,334.63
===================
Cheers,
Stan H.
===================